HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

PRE-BORN–OPPENHEIMER ENERGIES, LEADING-ORDER RELATIVISTIC and QED CORRECTIONS FOR ELECTRONICALLY

EXCITED STATES OF MOLECULAR HYDROGEN

\underline{E. SALY}, D. FERENC, E. MÁTYUS, ELTE, Eötvös Loránd University, Institute of Chemistry, Pázmány Péter sétány 1/A, Budapest, H-1117, Hungary

For rovibronic states corresponding to the B and B′ Σ_{u}^{+} electronic states of the hydrogen molecule, the pre-Born–Oppenheimer (four-particle) non-relativistic energy is converged to a 1–3 parts-per-billion relative precision. The four-particle nonrelativistic energy is appended with leading-order relativistic, leading- and estimated higher-order quantum-electrodynamics corrections [1]. The resulting term values referenced to the rovibronic ground state are obtained in an excellent agreement with the experimental results [2]. Further results are reported and discussed for other rovibronic states assignable to the C ^{1}Π_{u} and the EF, GK, and HH Σ_{g}^{+} electronic states.

References

[1] E. Saly, D. Ferenc, and E. Mátyus, Mol. Phys. e2163714 (2023).

[2] D. Bailly, E. J. Salumbides, M. Vervloet, and W. Ubachs, Mol. Phys. 108, 827 (2010).

p-number: p193  Submitted on Thu, 15 Jun 2023 16:32:51 +0200